A can of cat food, guaranteed by the manufacturer to contain at least 10 units of protein, 20 units of mineral matter, and 6 units of fat, consists of a mixture of four different ingredients. Ingredient A contains 10 units of protein, 2 units of mineral matter, and 1 2 unit of fat per 100g. Ingredient B contains 1 unit of protein, 40 units of mineral matter, and 3 units of fat per 100g. Ingredient C contains 1 unit of protein, 1 unit of mineral matter, and 6 units of fat per 100g. Ingredient D contains 5 units of protein, 10 units of mineral matter, and 3 units of fat per 100g. The cost of each ingredient is Birr 3, Birr 2, Birr 1, and Birr 4 per 100g, respectively. How many grams of each should be used to minimize the cost of the cat food, while still meeting the guaranteed composition? (Hint: Solve through simplex model)


1
Expert's answer
2022-01-03T07:18:12-0500

Solution:

"10A+B+C+5D\u226510,"

"2A+40B+C+10D\u226520,"

"12A+3B+6C+3D\u22656,"

"F = 3A+2B+C+4D~\rightarrow min,"

"\begin{cases}
 -10A-B-C-5D+K=-10\\
 -2A-40B-C-10D+L=-20\\
-12A-3B-6C-3D+M=-6
\end{cases}" "\def \arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c:c}
 C& 3 & 2&1&4&0&0&0&0 \\ \hline
 bas& A & B&C&D&K&L&M&b\\
 \hline
 K & -10 & -1&-1&-5&1&0&0&-10\\ \hline
L&-2&-40&-1&-10&0&1&0&-20\\ \hline
M&-12&-3&-6&-3&0&0&1&-6\\\hline
\end{array}" "|b|_{max}=|-20|"


"\def \arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c:c}
 C& 3 & 2&1&4&0&0&0&0 \\ \hline
 bas& A & B&C&D&K&L&M&b\\
 \hline
 K & -\frac{199}{20}& 0&-\frac{39}{40}&-\frac{19}{4}&1&-\frac{1}{40}&0&-\frac{19}{2}\\ \hline
B&\frac{1}{20}&1&\frac{1}{40}&\frac 14&0&-\frac{1}{40}&0&\frac 12\\ \hline
M&-\frac{237}{20}&0&-\frac{237}{40}&-\frac 94&0&-\frac{3}{40}&1&-\frac 92\\\hline
\end{array}" "|b|_{max}=|-\frac{19}{2}|"


"\def \arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c:c}
 C& 3 & 2&1&4&0&0&0&0 \\ \hline
 bas& A & B&C&D&K&L&M&b\\
 \hline
 A & 1& 0&\frac{39}{398}&\frac{95}{199}&-\frac{20}{199}&\frac{1}{398}&0&\frac{190}{199}\\ \hline
B&0&1&\frac{4}{199}&\frac {45}{199}&\frac{1}{199}&-\frac{5}{199}&0&\frac {90}{199}\\ \hline
M&0&0&-\frac{948}{199}&\frac {678}{199}&-\frac{237}{199}&-\frac{9}{199}&1&\frac {1356}{199}\\\hline
\end{array}"

"\def \arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c:c:c:c}
 C& 3 & 2&1&4&0&0&0&0 \\ \hline
 bas& A & B&C&D&K&L&M&b\\
 \hline
 A & 1& 0&\frac{39}{398}&\frac{95}{199}&-\frac{20}{199}&\frac{1}{398}&0&\frac{190}{199}\\ \hline
B&0&1&\frac{4}{199}&\frac {45}{199}&\frac{1}{199}&-\frac{5}{199}&0&\frac {90}{199}\\ \hline
M&0&0&-\frac{948}{199}&\frac {678}{199}&-\frac{237}{199}&-\frac{9}{199}&1&\frac {1356}{199}\\\hline
\Delta&0&0&-\frac{265}{398}&-\frac {421}{199}&-\frac{58}{199}&-\frac{17}{398}&0&\frac {750}{199}\\\hline
\end{array}"


"A=\frac{190}{199}, B =\frac{ 90}{199}, C = 0, D = 0,"

"F=3\cdot \frac{190}{199}+2\cdot \frac{90}{199}+1\cdot 0+4\cdot 0=3\frac{153}{199}\approx3.7688~\text{Birr}."